OpenStudy (jylarai):

Find the radius of a circle with an area of 100^2

OpenStudy (photon336):

\[Area = \pi*r^{2}\]

OpenStudy (photon336):

all you need to do is re-arrange your equation and solve for r. this is easy you know what the area is and you know what pi is. so to find the radius you just isolate r 1. by dividing both sides by pi 2. taking the square root of both sides. which gives you this formula below: \[\sqrt{\frac{ Area }{ \pi }} = r \]

OpenStudy (jylarai):

Ty very much for the help

OpenStudy (kayders1997):

So do you know what it is? :)

OpenStudy (kayders1997):

And does it says 100pi?

OpenStudy (mathmale):

Double check the original problem. Does it say that the area of the circle is 100 pi, or does it say 100^2?